Brian is three times as old as Jrodan. maggie is five years older than Jordan. If the sum of their ages is 100, how old are Brian, Jordan, and Maggie
Answer by ankor@dixie-net.com(22740)

Brian is three times as old as Jordan.
b = 3j
maggie is five years older than Jordan.
m = j+5
If the sum of their ages is 100,
b + j + m = 100
how old are Brian, Jordan, and Maggie
In the above equation, replace b with 3j and replace m with (j+5)
3j + j + (j+5) = 100
5j = 100 - 5
5j = 95
j = 95/5
j = 19 is Jordan's age
:
Use the first two equations to find b and m. See that your 3 solutions add up to 100